# NeuMaDiff: Neural Material Synthesis via Hyperdiffusion

### v1

@@ 3.3 Multi-Modal Conditional Hyperdiffusion
To model the complex distribution within the NeuMERL dataset, we utilize a diffusion process [25, 41, 13]. Specifically, we employ a transformer-based denoising network [52], leveraging its demonstrated efficacy [41] and its attention mechanisms allowing for an effective focus on relevant information, enhancing the network’s ability to capture intricate dependencies in the data.
Our hyperdiffusion supports conditioning across three modalities: material type (represented as integers), text description, and reference images. For each modality, we utilize a categorical encoding for material type, an augmented CLIP text embedding [43, 62] for text, and a ResNet [22] for images. This multi-modal conditioning approach enables material synthesis to be guided by different user inputs, enhancing workflow intuitiveness and accessibility and allowing for a more smooth and accurate translation of creative vision into generated materials. For conditional sampling, we employ classifier-free guidance (CFG) [24].
Please refer to Secs. A.2, B.2, A.3 and B.3 in the supplementary for further details on the diffusion model, transformer, attention mechanism, and CFG, respectively.
@@ 4 Experiments
Figure 6: Material synthesis across various pipelines. All baseline models fail to capture the underlying distribution effectively, resulting in meaningless samples with severe artefacts in the synthesized materials. In contrast, NeuMaDiff successfully captures the complex neural material distribution, achieving significantly better fidelity and diversity.
In this section, we present an extensive evaluation of NeuMaDiff, beginning with an overview of our datasets (Sec. 4.1) and introducing novel material distributional metrics (Sec. 4.2). We then demonstrate the effectiveness of our pipeline through experiments on unconditional synthesis (Sec. 4.3), multi-modal conditional synthesis (Sec. 4.4), and constrained synthesis (Sec. 4.5).
For additional details, refer to the supplementary material: model implementation and training specifics are discussed in Secs. B and C.1, further results are presented in Sec. E, and Sec. F includes additional experiments.
@@ 4.1 Dataset
We fit neural fields to individual materials in the AugMERL dataset (Sec. 3.1), which is derived from the MERL BRDF dataset [34]. Our hyperdiffusion model is trained on the NeuMERL dataset (Sec. 3.2), which consists of neural material representations. The training-validation split is 80%-20% for AugMERL and 95%-5% for NeuMERL. In the constrained synthesis experiments in Sec. 4.5, statistical information is gathered from the MERL dataset [34].
@@ 4.2 Material Distributional Metrics
We use Fréchet Inception Distance (FID) [23] as an image-based metric to assess the quality of rendered single-view images by comparing them to a reference set.
To the best of our knowledge, effective metrics for directly comparing material distributions are still lacking. Drawing inspiration from distributional metrics for point clouds [2, 32, 59], we introduce three novel material distributional metrics – minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor (1-NNA) – to evaluate the fidelity and diversity of synthesized BRDF sets $\mathcal{S}$ relative to a reference set $\mathcal{R}$. Each metric is based on an underlying distance measure $d(f_{r},f_{r}^{\prime})$ between two BRDFs.
@@
$\mathcal{L}_{\text{MMD}}^{d}(\mathcal{R},\mathcal{S}):=\frac{1}{\left|\mathcal{R}\right|}\sum_{f_{r}\in\mathcal{R}}\underset{f_{r}^{\prime}\in\mathcal{S}}{\min}d(f_{r},f_{r}^{\prime})$ (3)
@@
$\mathcal{L}_{\text{1-NNA}}^{d}(\mathcal{R},\mathcal{S}):=\frac{\sum\limits_{f_{r}\in\mathcal{R}}\mathbb{I}\left[N_{f_{r}}\in\mathcal{R}\right]+\sum\limits_{f_{r}^{\prime}\in\mathcal{S}}\mathbb{I}[N_{f_{r}^{\prime}}\in\mathcal{S}]}{\left|\mathcal{S}\right|+\left|\mathcal{R}\right|},$ (5)
where $\mathbb{I}[\cdot]$ is the indicator function and $N_{f_{r}}$ denotes the nearest neighbor of $f_{r}$ in $(\mathcal{R}\cup\mathcal{S})-\{f_{r}\}$. In this metric, each sample is classified as belonging to either the reference set $\mathcal{R}$ or the synthesized set $\mathcal{S}$ based on the membership of its nearest neighbor. If $\mathcal{R}$ and $\mathcal{S}$ are drawn from the same underlying distribution, the classifier’s accuracy will approach 50% with a large sample size. An accuracy closer to 50% indicates greater similarity between $\mathcal{R}$ and $\mathcal{S}$, suggesting that the model has effectively learned the target distribution.
Each distributional metric can be computed using an underlying distance $d$. Potential options include rendering-based metrics such as root mean squared error (RMSE), peak signal-to-noise ratio (PSNR), and structural similarity index measure (SSIM) [54]. PSNR measures the reconstruction quality of images, while SSIM predicts perceived similarity. Since higher values in PSNR and SSIM correspond to greater similarity, we negate these values – resulting in NegPSNR and NegSSIM, respectively – to make them plausible distance functions [^1] Although NegPSNR and NegSSIM may not strictly meet the criteria for distance metrics (e.g., violating positivity), they suffice to provide useful relative distance measures for material distributional metrics.. To directly assess the distance between two BRDFs without relying on renderings, we also introduce the following BRDF-L1 distance:
$d_{\text{BRDF-L1}}:=\underset{\theta_{{\bm{\mathbf{h}}}},\theta_{{\bm{\mathbf{d}}}},\varphi_{{\bm{\mathbf{d}}}}}{\mathbb{E}}\Big{[}\left|f_{r}-f_{r}^{\prime}\right|\Big{]};$ (6)
For further background on the image-based metrics and validation of the proposed material distributional metrics, please refer to Sec. D in the supplementary.
@@ 4.3 Unconditional Synthesis
We begin by presenting the results of unconditional synthesis of neural materials using NeuMaDiff. Most existing approaches utilize principal component analysis (PCA) [1, 48] and variational autoencoders (VAEs) [28, 17]. Accordingly, we developed four generative baselines based on the two methods PCA and VAE. Each method was applied to either the AugMERL or NeuMERL datasets, resulting in the following baselines: PCA-A, PCA-N, VAE-A, and VAE-N. For synthesis, PCA-based baselines perform random linear interpolation in the principal space, while VAE-based baselines sample from a standard Gaussian prior. Additional background details on these methods can be found in Secs. A.4 and B.4 in the supplementary material.
Table 1 provides a detailed comparison of synthesized materials using our proposed metrics, while Figs. 6, 5 and 1 showcase renderings of these materials across various scenes, from simple spheres to more complex models. The quantitative and qualitative results indicate that NeuMaDiff consistently outperforms all baselines across metrics, producing diverse, high-quality, visually appealing, and perceptually realistic renderings. This demonstrates the effectiveness of NeuMaDiff for neural material synthesis. Notably, materials synthesized by the PCA- and VAE-based baselines exhibit significant artifacts, likely due to the limitations of these simpler models in capturing the complex distribution of neural materials.
Metric Training set PCA-A PCA-N VAE-A VAE-N NeuMaDiff (ours)
FID ($\downarrow$) 0.187 10.9 23.8 26.1 28.4 0.440
MMD ($\downarrow$) BRDF-L1×$\times 10^{-3}$ 2.51 9.05 9.22 9.09 $\infty$ 4.02
RMSE×$\times 10^{2}$ 7.54 33.3 30.2 63.7 69.2 9.34
NegPSNR --28.7 --13.9 --14.8 --8.30 --7.40 --25.6
NegSSIM×$\times 10$-9.55 --6.74 --6.29 --2.68 --2.15 --9.40
COV (%) ($\uparrow$) BRDF-L1 60.8 2.50 30 0.833 2.50 50.8
RMSE 55.8 18.3 28.3 0.833 0.833 50.0
NegPSNR 56.7 18.3 28.3 0.833 0.833 50.0
NegSSIM 59.2 23.3 16.7 0.833 0.833 51.7
1-NNA (%) ($\downarrow$) BRDF-L1 58.8 100 95.4 100 100 80.0
RMSE 55.4 96.3 93.4 100 100 60.0
NegPSNR 55.0 94.2 90.0 100 100 60.4
NegSSIM 57.5 96.3 96.7 100 100 61.7
Table 1: Qualitative evaluation of unconditional synthesis with metrics assessing generation fidelity and diversity. $\downarrow$ indicates that a lower score is better and $\uparrow$ indicates the opposite. NeuMaDiff significantly outperforms all baseline models across these metrics, underscoring its effectiveness in neural material synthesis.
@@ 4.4 Multi-Modal Conditional Synthesis
For material type conditioning, we represent each of the 48 material types in the MERL dataset [34] (e.g., acrylic, aluminium, metallic, plastic, etc.) using integers. The full list of material types is available in Sec. C.2 in the supplementary; For text-based conditioning, we use descriptions derived from the MERL dataset materials [34]. For additional materials in the AugMERL dataset, descriptions are assigned as follows: For RGB-permuted materials, we retain the original description but omit color-specific words (e.g., “red metallic paint” becomes “metallic paint”); For PCA-interpolated materials, we generate descriptions in the format “a mixture of $t_{A}$ and $t_{B}$, where $t_{A}$ and $t_{B}$ are the descriptions of the interpolated materials $A$ and $B$, respectively; For image-based conditioning, we use cropped single-view renderings of materials from AugMERL as input to guide the synthesis based on visual references.
Figures 8, 9 and 10 present the results for type-, text-, and image-conditioned synthesis, respectively. Across all conditioning modes, the synthesized materials demonstrate realism, diversity, and a close alignment with the input conditions. Notably, in text-conditioned synthesis, NeuMaDiff effectively generalizes to novel inputs, such as “green metal”, “red plastic”, and “highly specular material”, producing materials that are perceptually consistent with these previously unseen descriptions.
Figure 8: Type-conditioned synthesis. The synthesized materials are diverse and closely align with the specified material type.
Figure 9: Text-conditioned synthesis. NeuMaDiff synthesizes materials aligning with the texts and generalizes to unseen inputs: “green metal”, “red plastic”, and “highly specular material”.
Figure 10: Image-conditioned synthesis. In each of the six pairs, the left image is the conditioning input, while the right image is the synthesized result. NeuMaDiff effectively generates realistic materials that closely align with the conditioning images.
@@ 4.5 Constrained Synthesis
We derive the theoretical upper limit for a diffuse reflectance value, $f_{\text{diffuse}}$, for a constant (i.e., purely diffuse) BRDF $f_{r}({\bm{\mathbf{\omega}}}_{i},{\bm{\mathbf{\omega}}}_{o})\equiv f_{\text{diffuse}}$ in each color channel. For a physically valid BRDF that adheres to energy passivity [63], the reflected energy must not exceed the incident energy in each channel. Thus, we have:
11 11 $\geq\int_{H^{2}}f_{r}({\bm{\mathbf{\omega}}}_{i},{\bm{\mathbf{\omega}}}_{o})\cos\theta_{o}\,\mathrm{d}{\bm{\mathbf{\omega}}}_{o}$ (7)
$=f_{\text{diffuse}}\int_{H^{2}}\cos\theta_{o}\,\mathrm{d}{\bm{\mathbf{\omega}}}_{o}=\pi f_{\text{diffuse}}$ (8)
$\Rightarrow f_{\text{diffuse}}\leq\frac{1}{\pi}.$ (9)
Building on this observation and a statistical analysis of the MERL’s mean and maximum reflectance values across color channels and material types (Sec. F.1 in the supplementary), we propose a novel set of rules for categorizing materials into seven types. These rules enable the constrained selection of synthesized materials based on desired material characteristics. Unlike many black-box machine learning approaches, this method is rooted in BRDF analysis, offering inherent explainability and interpretability. Below, we outline two of these rules, with the full set detailed in Sec. C.3.
  • Purely diffuse materials: The reflectance values in all directions do not exceed the diffuse threshold $f_{\text{diffuse}}$, allowing for only $e:=8\times 10^{4}$ exceptions:
    |$\lvert\{({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\in\mathbb{R}^{6}\mid\left\|f_{r}({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\right\|_{\infty}>f_{\text{diffuse}}\}\rvert<e,$ (10)
@@
Figure 7 shows materials generated through our constrained synthesis, incorporating the proposed filtering rules to ensure that the synthesized outputs match the specified material categories. The results demonstrate that the synthesized materials effectively exhibit the characteristics of the desired categories.
5 Conclusion and Discussion
High-quality material synthesis is critical for achieving realistic, convincing renderings. In this work, we introduced NeuMaDiff, an innovative neural material synthesis pipeline built upon a multi-modal conditional hyperdiffusion model. Using neural fields as the core representation for materials, we trained our hyperdiffusion on their weights, demonstrating its effectiveness in generating high-quality, diverse materials through extensive experiments. Additionally, we contribute two material datasets and two BRDF distributional metrics to facilitate future research.
However, our current evaluation is limited to uniform material datasets. Extending our approach to accommodate more complex materials, such as spatially varying BRDFs, may require additional mechanisms, which we aim to explore in future work.
Supplementary Material
A Additional Background
A.1 Principal Component Analysis
Principal component analysis (PCA) [1], also known as Karhunen-Loève transform or Hotelling transform, is a linear dimensionality reduction technique. We utilize PCA for data augmentation in Sec. A.1 and baseline models in Sec. 4.3.

### v2

@@ 3.3 Multi-Modal Conditional Hyperdiffusion
To model the complex distribution within the NeuMERL dataset, we utilize a diffusion process (ho2020denoisingdiffusion; peebles2022learning; erkoc2023hyperdiffusion). Specifically, we employ a transformer-based denoising network (vaswani2017attention), leveraging its demonstrated efficacy (peebles2022learning) and its attention mechanisms allowing for an effective focus on relevant information, enhancing the network’s ability to capture intricate dependencies in the data.
Our hyperdiffusion supports conditioning across three modalities: material type (represented as integers), text description, and reference images. We utilize a categorical encoding for material type, an augmented CLIP text embedding (radford2021learning; zhou2023clip) for text, and a ResNet (he2016deep) for images. This multi-modal conditioning approach enables material synthesis to be guided by different user inputs, enhancing workflow intuitiveness and accessibility and allowing for a more smooth and accurate translation of creative vision into generated materials. For conditional sampling, we employ classifier-free guidance (CFG) (ho2022cfg).
Please refer to Secs.˜B.2, C.2, B.3 and D.2 in the supplementary for further details on the diffusion model, transformer, attention mechanism, and CFG, respectively.
@@ 4 Experiments
Figure 5: Material synthesis. Baseline models fail to capture the underlying distribution effectively, resulting in homogeneous outputs or severe artifacts. In contrast, M3ashy successfully captures the complex neural material distribution, achieving significantly better fidelity and diversity. Our materials also support spatially varying rendering configurations (last three columns).
In this section, we present extensive experiments on M3ashy. For additional details, please refer to the appendix: model and experiment specifics in Appendices˜C and D, further results in Appendix˜F, and more experiments in Appendix˜G.
@@ 4.1 Dataset
We fit neural fields to individual materials in the AugMERL dataset (Sec.˜3.1), which is derived from the MERL BRDF dataset (Matusik2003datadriven). Our hyperdiffusion model is trained on the NeuMERL dataset (Sec.˜3.2), which consists of neural material representations. The training-validation split is 80%-20% for each individual fitting on AugMERL and 95%-5% for NeuMERL. All samples derived from a given material remain within one split. In the constrained synthesis experiments in Sec.˜4.5, statistical information is gathered from the MERL dataset (Matusik2003datadriven).
@@ 4.2 Material Distributional Metrics
We use Fréchet Inception Distance (FID) (heusel2017gans) as an image-based metric to assess the quality of rendered single-view images.
To the best of our knowledge, effective metrics directly comparing material distributions are still lacking. Drawing inspiration from the metrics for point clouds (yang2019pointflow), we introduce three novel material distributional metrics – minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor (1-NNA) – to evaluate the fidelity and diversity of synthesized BRDF sets $\mathcal{S}$ relative to a reference set $\mathcal{R}$. Each metric is based on an underlying distance measure $d(f_{r},f_{r}^{\prime})$ between two BRDFs.
@@
$\mathcal{L}_{\text{MMD}}^{d}(\mathcal{R},\mathcal{S})\coloneqq\frac{1}{\left|\mathcal{R}\right|}\sum_{f_{r}\in\mathcal{R}}\underset{f_{r}^{\prime}\in\mathcal{S}}{\min}d(f_{r},f_{r}^{\prime})$ (3)
@@
$\mathcal{L}_{\text{1-NNA}}^{d}(\mathcal{R},\mathcal{S})\coloneqq\frac{\sum\limits_{f_{r}\in\mathcal{R}}\mathbb{I}\left[N_{f_{r}}\in\mathcal{R}\right]+\sum\limits_{f_{r}^{\prime}\in\mathcal{S}}\mathbb{I}[N_{f_{r}^{\prime}}\in\mathcal{S}]}{\left|\mathcal{S}\right|+\left|\mathcal{R}\right|},$ (5)
where $\mathbb{I}[\cdot]$ is the indicator function and $N_{f_{r}}$ denotes the nearest neighbor of $f_{r}$ in $(\mathcal{R}\cup\mathcal{S})-\{f_{r}\}$. In this metric, each sample is classified as belonging to either the reference set $\mathcal{R}$ or the synthesized set $\mathcal{S}$ based on the membership of its nearest neighbor. If $\mathcal{R}$ and $\mathcal{S}$ are drawn from the same underlying distribution, the classifier’s accuracy will approach 50% with a large sample size.
Each metric can be computed using an underlying distance $d$. Potential options include rendering-based metrics such as root mean squared error (RMSE), peak signal-to-noise ratio (PSNR), and structural similarity index measure (SSIM) (SSIM04). Since higher values in PSNR and SSIM correspond to greater similarity, we negate these values – resulting in NegPSNR and NegSSIM, respectively – to make them plausible distance functions To directly assess the distance between two BRDFs without relying on renderings, we also introduce the following BRDF L1 distance:
$d_{\text{BRDF-L1}}\coloneqq\underset{\theta_{{\bm{\mathbf{h}}}},\theta_{{\bm{\mathbf{d}}}},\varphi_{{\bm{\mathbf{d}}}}}{\mathbb{E}}\Big[\left|f_{r}-f_{r}^{\prime}\right|\Big];$ (6)
For further background on the image-based metrics and validation of the proposed material distributional metrics, please refer to Appendix˜E in the supplementary.
@@ 4.3 Unconditional Synthesis
We begin by presenting the results of unconditional synthesis of neural materials using M3ashy. We compare our approach with the PCA-based method of NielsenPCA2015 and with the sparse reconstruction model of gokbudak2023hypernetworks, which uses a hypernetwork to model measured tabular BRDF data. As gokbudak2023hypernetworks’s method is not generative, we extend it with a variational autoencoder (VAE) to enable comparison. Both methods are applied to the AugMERL (-A) and NeuMERL (-N) datasets, resulting in four baselines: VAE-A, VAE-N, PCA-A, and PCA-N. We also develop another baseline, MERL100, that represents our method trained on the original MERL dataset to demonstrate the effectiveness of our data augmentation.
Table˜2 provides a detailed comparison of baselines with our proposed metrics, while Figs.˜5 and 1 showcase renderings of these materials across various geometries and scenes. To achieve more complex visual effects, our synthesized materials support rendering with bump or normal maps, as well as spatially varying configurations (Jakob2022DrJit) (additional results in Appendix˜F). The quantitative and qualitative results indicate that M3ashy consistently outperforms all baselines across metrics, producing diverse, high-quality, visually appealing, and perceptually realistic renderings. This demonstrates the effectiveness of M3ashy for neural material synthesis. Notably, materials synthesized by some baselines exhibit significant artifacts, likely due to the limitations of these simpler models in capturing the complex distribution of measured materials.
Metric Training set PCA-A PCA-N VAE-A VAE-N MERL100 M3ashy (ours)
FID ($\downarrow$) 0.187 10.9 23.8 26.1 10.0 7.56 0.440
[…]
NegPSNR -28.7 -13.9 -14.8 -8.30 -20.9 -22.6 -25.6
[…]
COV (%) ($\uparrow$) BRDF-L1 60.8 2.50 30 0.833 20.8 28.3 50.8
RMSE 55.8 18.3 28.3 0.833 16.7 25.0 50.0
NegPSNR 56.7 18.3 28.3 0.833 18.3 25.0 50.0
NegSSIM 59.2 23.3 16.7 0.833 17.5 22.5 51.7
1-NNA (%) ($\downarrow$) BRDF-L1 58.8 100 95.4 100 96.7 92.5 80.0
RMSE 55.4 96.3 93.4 100 93.3 84.6 60.0
NegPSNR 55.0 94.2 90.0 100 93.3 84.6 60.4
NegSSIM 57.5 96.3 96.7 100 93.8 86.3 61.7
Table 2: Qualitative evaluation of unconditional synthesis with metrics assessing generation fidelity and diversity. M3ashy significantly outperforms all baseline models across these metrics, underscoring its effectiveness in neural material synthesis.
@@ 4.4 Multi-Modal Conditional Synthesis
For material type conditioning, we represent each of the 48 material types in the MERL dataset (Matusik2003datadriven) (e.g., acrylic, metallic, plastic, etc.) using integers. The full list of material types is available in Sec.˜D.3 in the supplementary; For text conditioning, we use descriptions derived from the MERL dataset (Matusik2003datadriven). For additional materials in the AugMERL dataset, descriptions are assigned as follows: For RGB-permuted materials, we retain the original description but omit color-specific words (e.g., “red metallic paint” becomes “metallic paint”); For PCA-interpolated materials, we generate descriptions in the format “a mixture of $t_{A}$ and $t_{B}$, where $t_{A}$ and $t_{B}$ are the descriptions of the interpolated materials $A$ and $B$, respectively; For image-based conditioning, we use cropped single-view renderings of materials from AugMERL as input. We encode input texts and images using CLIP encoders (radford2021learning).
Figures˜7, 8 and 9 present the results for type-, text-, and image-conditioned synthesis, respectively. Across all conditioning modes, the synthesized materials demonstrate realism, diversity, and a close alignment with the input conditions. Notably, in text- and image-conditioned synthesis, M3ashy effectively generalizes to unseen texts (e.g., “green metal”, “red plastic”, and “highly specular material”) and real-world images, producing materials that are perceptually consistent with these previously unseen inputs.
Figure 7: Type-conditioned synthesis. The synthesized materials are diverse and closely align with the input type.
Figure 8: Text-conditioned synthesis. Synthesized materials align with the texts and generalize to unseen inputs: “green metal”, “red plastic”, and “highly specular material”.
Figure 9: Image-conditioned synthesis. Each of the eight pairs consists of the input (left) and the synthesized (right). M3ashy effectively generates realistic materials that closely align with the conditioning images and generalizes to unseen, real-world images (last column).
@@ 4.5 Constrained Synthesis
We derive the theoretical upper limit for a diffuse reflectance value, $f_{\text{diffuse}}$, for a constant (i.e., purely diffuse) BRDF $f_{r}({\bm{\mathbf{\omega}}}_{i},{\bm{\mathbf{\omega}}}_{o})\equiv f_{\text{diffuse}}$ in each color channel. For a physically valid BRDF that adheres to energy passivity (zhou2024physically), the reflected energy must not exceed the incident energy in each channel. Thus, we have:
$1$ $\geq\int_{H^{2}}f_{r}({\bm{\mathbf{\omega}}}_{i},{\bm{\mathbf{\omega}}}_{o})\cos\theta_{o}\,\mathrm{d}{\bm{\mathbf{\omega}}}_{o}$ (7)
$=f_{\text{diffuse}}\int_{H^{2}}\cos\theta_{o}\,\mathrm{d}{\bm{\mathbf{\omega}}}_{o}=\pi f_{\text{diffuse}}\Rightarrow f_{\text{diffuse}}\leq\frac{1}{\pi}.$ (8)
Building on this observation and a statistical analysis of the MERL’s mean and maximum reflectance values across color channels and material types (Sec.˜G.1 in the supplementary), we propose a set of rules for categorizing materials into seven types. These rules enable the selection of synthesized materials based on desired material characteristics. Unlike many black-box machine learning approaches, this method is rooted in BRDF analysis, offering inherent explainability and interpretability. Below, we outline two of these rules, with the full set detailed in Sec.˜D.4.
  • Purely diffuse materials: The reflectance values in all directions do not exceed the diffuse threshold $f_{\text{diffuse}}$, allowing for only $e\coloneqq 8\times 10^{4}$ exceptions:
    $\lvert\{({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\in\mathbb{R}^{6}\mid\left\|f_{r}({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\right\|_{\infty}>f_{\text{diffuse}}\}\rvert<e,$ (9)
@@
Figure˜6 shows materials generated through our constrained synthesis, incorporating the proposed filtering rules to ensure that the synthesized outputs match the specified material categories. The results demonstrate that the synthesized materials effectively exhibit the characteristics of the desired categories.
4.6 Ablation Study
To assess the impact of our augmented material dataset AugMERL, we further train our model on the original MERL dataset in the unconditional synthesis task. We report both quantitative and qualitative results for this model, presented in the “MERL100” columns of Tab.˜2 and Fig.˜5, respectively. The results indicate that the model trained on AugMERL exhibits higher quality and diversity compared to the one trained on MERL, demonstrating the effectiveness of our augmented dataset in enhancing the synthesis pipeline.
Additionally, we conduct sparse BRDF reconstruction and BRDF compression experiments following a previous method (gokbudak2023hypernetworks). For sparse reconstruction, we set the sample size to $N=4000$, while for compression, we use a latent dimension of $40$. In both experiments, we train the model on either the original MERL dataset or the AugMERL dataset. The results, summarized in Tab.˜3, demonstrate that training on AugMERL consistently enhances material quality across all evaluated metrics, further validating the effectiveness of our augmented dataset.
Metric Sparse reconstruction Compression
MERL AugMERL MERL AugMERL
PSNR ($\uparrow$) 32.2 36.3 45.2 48.3
Delta E ($\downarrow$) 2.1 1.8 0.693 0.623
SSIM ($\uparrow$) 0.972 0.983 0.994 0.994
Table 3: Quantitative comparison of training on MERL versus AugMERL in the sparse BRDF reconstruction and BRDF compression experiments. The results demonstrate that training on AugMERL consistently enhances performance across all metrics.
5 Conclusion and Future Work
We introduced M3ashy, a multi-modal material synthesis approach with hyperdiffusion. Using neural fields as the core representation, we trained hyperdiffusion on their weights, demonstrating its ability to generate high-quality, diverse materials. Additionally, we contribute two material datasets and three BRDF metrics for future research. At this stage, our method does not account for physical correctness. Promising directions for future work include developing physically accurate neural representations of BRDFs and extending the approach to support more complex materials.
Appendix A Related Work on Material Acquisition and Databases
Capturing real-world material appearance requires lengthy acquisition times and substantial storage. Traditional capture uses four-axis gonioreflectometers (gonio; White98) with later advancements enhancing angular coverage, wavelength resolution, efficiency, and accuracy (White98). Systems range from basic sensor setups, like camera arrays (weinmann2015advances), varied illumination sources (haindl2013visual), and specific geometries (marschner99; kaleidoscop; weinmann2015material), to advanced configurations for capturing full material patches (danabtf; weinmann-2014). Filip and Vávra contributed a database of 150 materials, many exhibiting anisotropic properties (filip2014template). These material databases are continuously being augmented by later methods (ngan2005experimental; dupuy2018adaptive). These improvements have enabled the creation of extensive BRDF material databases (Matusik2003datadriven; rgl2018; NielsenPCA2015), which underpin the datasets in our work.
Appendix B Additional Background
B.1 Principal Component Analysis
Principal component analysis (PCA) (abdi2010principal), also known as Karhunen-Loève transform or Hotelling transform, is a linear dimensionality reduction technique. We utilize PCA for data augmentation in Sec.˜B.1 and baseline models in Sec.˜4.3.